in a factory, 6% of the workers wear glasses. Suppose 20 workers are randomly selected. Find the probability of two workers wearing glasses, no worker wearing glasses, three or more workers wearing glasses
Answer by Boreal(15235)

for 2, 20C2*0.06^2*0.94^18=0.2246
for 0, it is 0.94^20=0.2901
for 1 it is 20*0.06*0.94^19=03703
The probability for those 3 outcomes is 0.8850
The probability for 3 or more is the complement or 0.1150.